# Federated Unsupervised Domain Generalization using Global and Local Alignment of Gradients

### v1

@@ 3.3 FedGaLA
The analysis above establishes a link between gradient alignment and domain shift, forming the basis for our proposed FedGaLA framework. Figure 1 illustrates an overview of our framework for addressing the newly proposed problem setup (Definition 1). FedGaLA is remotely inspired by prior works that demonstrate improved generalization in centralized (non-federated) learning through gradient alignment [29, 42, 43]. The study in [29] updates the weights if the signs of the gradient components are aligned across all domains. Fish [42] maximizes the inner product between gradients from different domains, while Fishr [43] leverages domain-level gradient variances. However, our framework extends this notion by integrating gradient alignment into the field of unsupervised and federated learning, and assumes, unlike [29, 42, 43], that the distribution of data from different clients is not known. Our core idea includes (i) enabling clients to learn domain-invariant representations at the client level through local gradient alignment, and (ii) adjusting the aggregation weights at the server level using global gradient alignment. At each communication round, clients are initialized with the global model. Subsequently, each client updates its parameters using SSL for $E$ epochs based on the local data through local gradient alignment and sends these updates back to the server. Finally, the server employs the global gradient alignment technique to perform aggregation. This procedure is repeated for $T$ communications rounds to determine the global model. The remaining part of this section provides details on local and global gradient alignment, and the complete framework is outlined in Algorithm 1 (Appendix B).
@@
Local gradient alignment. Our method performs layer-wise local gradient alignment using a reference gradient. This reference is derived from the $l^{\text{th}}$ layer of the global model’s parameter updates between the current and the previous communication round. Suppose $\Theta=\{\theta^{(l)}\}_{l=1}^{L}$ indicates the parameters of the global model, where $\theta^{(l)}$ represents the parameters of the $l^{\text{th}}$ layer. The reference gradient for the $l^{th}$ layer is computed as $\hat{\textbf{g}}_{est}^{(l,t)}=\theta^{(l,t)}-\theta^{(l,t-1)}$, where $\theta^{(l,t)}$ and $\theta^{(l,t-1)}$ are the parameters of the $l^{\text{th}}$ layer of the global model at rounds $t$ and $t-1$, respectively. Then, $\hat{\textbf{g}}_{est}^{(l,t)}$ is locally used to determine whether the gradient of each layer obtained during training (e.g., via SGD) is aligned with the reference. The cosine similarity between the batch gradient and the reference for each layer $l$ at round $t$ is computed as $cos(\phi)^{(l,t)}=\frac{\langle\textbf{g}_{i,k}^{(l,t)},\textbf{g}_{\text{est}}^{(l,t)}\rangle}{\|\textbf{g}_{i,k}^{(l,t)}\|\cdot\|\textbf{g}_{\text{est}}^{(l,t)}\|}$, where $\textbf{g}_{i,k}^{(l,t)}$ is the gradient of $k^{\text{th}}$ batch of the $i^{\text{th}}$ client at layer $l$ and round $t$. Finally, batch gradients whose similarity with $\hat{\textbf{g}}_{est}^{(l,t)}$ are less than a user-defined threshold $\tau$ are discarded during the update process. This prevents clients from learning domain-specific features by disregarding local gradients that are not aligned with the global model.
@@
Global gradient alignment. To aggregate the local (client) models at the server, the locally measured gradients that closely match the average gradient across all clients are assigned greater weights. The process operates as follows. Once the server receives the local models, it first obtains the local update $\hat{\textbf{g}}_{i}^{(t)}=\Theta_{i}^{(t)}-\Theta^{(t-1)}$. Then, the initial global update $\hat{\textbf{g}}^{(t+1)}$ is calculated by averaging all $\hat{\textbf{g}}_{i}^{(t)}$. Subsequently, the weight of each client is computed using $w_{i}=\frac{cos(\hat{\textbf{g}}_{i}^{(t)},\hat{\textbf{g}}^{(t+1)})+1}{2}$, where $w_{i}$ is the weight for the $i^{\text{th}}$ client and $\hat{\textbf{g}}_{i}^{(t)}$ is the gradient of the $i^{\text{th}}$ client at round $t$. This weight reflects the degree of alignment between each client’s update and the global model’s update direction. To ensure these weights are properly normalized using $w_{i}={w_{i}}/{\sum_{k=1}^{K}w_{k}}$. The normalization of weights allows for the proportional contribution of each client’s update.
@@
Finally, the normalized weights are used to perform aggregation. Each client’s model update is scaled by its respective weight, and these weighted updates are then aggregated to compute the weighted average update. The global model at round $t+1$ is updated based on $\hat{\textbf{g}}^{(t+1)}=\frac{1}{K}\sum_{i=1}^{K}w_{i}\hat{\textbf{g}}_{i}^{(t)}$. This aggregation step is repeated three times, refining the weights with each iteration. This ensures the global model update is significantly influenced by clients whose updates align closely with the global learning objective. After completing the aggregation process, the global model is further updated based on $\Theta^{(t+1)}\leftarrow\Theta^{(t)}+\hat{\textbf{g}}^{(t+1)}$.
@@ 4 Experiment setup
Datasets. To evaluate the effectiveness of our proposed method, we conduct experiments across four commonly used benchmarks for domain generalization. They include: PACS [44], which consists of 9,991 images from four domains: ‘Photo’, ‘Art-painting’, ‘Cartoon’, and ‘Sketch’, across seven classes; Office-Home [13], which consists of 15,588 images from four domains: ‘Art’, ‘Clipart’, ‘Product’, and ‘Real-world’, across 65 classes; TerraInc [15], which includes 24,788 images from four domains ‘Location 38’, ‘Location 43’, ‘Location 46’, and ‘Location 100’, across nine classes; DomainNet [14], which consists of 569,010 images in six domains: ‘Clipart’, ‘Infograph’, ‘Painting’, ‘Quickdraw’, ‘Real’, and ‘Sketch’, covering 345 classes. We follow the same protocol as [45] for this dataset (see details in Appendix D.1).
Evaluation. We use the leave-one-domain-out setting used in prior works [28, 45, 46]. This involves selecting one domain as the target, training the model on the rest of the domains, and then testing the model’s performance on the selected target domain. Linear evaluation, a common feature evaluation approach, is utilized to evaluate the quality of learned representations [47, 48, 49]. For linear evaluation, following [26, 27], we utilize 10% and 30% of the target data to train the linear classifier, and evaluate the remaining 90% and 70% of the data, respectively.
Model Labeled PACS DomainNet ratio P A C S Ave. C I Q Ave. FedEMA 10% 50.0(0.7) 29.5(1.9) 42.4(2.3) 45.6(0.16) 41.9 38.6(1.1) 13.7(0.5) 45.5(1.6) 32.4 FedBYOL 52.1(1.1) 31.8(1.1) 45.4(2.2) 47.4(1.9) 44.2 38.1(0.5) 14.1(0.4) 53.6(2.9) 31.8 FedMoCo 58.5(2.2) 35.7(9.6) 37.7(12.9) 36.6(8.2) 42.1 30.5(0.6) 10.9(2.1) 46.4(0.8) 27.2 FedSimSiam 46.2(1.1) 28.6(1.0) 46.7(0.6) 37.6(1.3) 39.8 44.8(1.5) 12.2(0.3) 40.3(2.4) 36.9 FedSimCLR 64.2(1.2) 41.9(1.5) 58.4(1.3) 70.1(1.2) 58.6 45.2(0.4) 13.7(0.3) 59.7(0.7) 39.5 FedGaLA (ours) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0 47.6(0.9) 14.2(0.5) 61.4(0.3) 41.1 FedEMA 30% 33.9(2.8) 48.3(2.4) 53.5(0.3) 45.3(2.3) 45.2 43.5(0.8) 20.0(1.5) 49.0(2.8) 37.5 FedBYOL 55.1(1.2) 35.3(1.5) 48.3(1.5) 48.9(0.5) 46.9 44.6(1.2) 20.5(1.1) 50.4(0.9) 38.5 FedSimSiam 58.5(2.2) 35.7(9.6) 37.7(12.9) 36.6(8.2) 42.1 35.7(1.1) 14.9(3.9) 44.6(3.5) 31.8 FedMoCo 47.3(1.6) 30.2(3.4) 47.4(0.7) 34.4(1.7) 39.9 51.5(1.8) 18.2(0.1) 59.9(3.6) 43.2 FedSimCLR 69.8(1.1) 46.4(2.1) 63.9(1.6) 73.4(3.0) 63.3 51.7(1.0) 16.3(0.2) 66.5(0.9) 44.8 FedGaLA (ours) 71.1(2.0) 46.8(2.1) 65.7(1.6) 74.5(1.1) 64.6 52.4(0.7) 16.2(0.6) 68.8(0.6) 45.8
@@
Baselines. To evaluate our method, we take a two-pronged approach: (1) We adapt several popular SSL approaches to the federated domain generalization task, denoting them as FedSimCLR, FedMoCo, FedBYOL, and FedSimSiam. To do so, we first train each client locally using the respective SSL method. Next, we aggregate the trained encoders at the server using FedAVG [1]. For BYOL and SimSiam, we follow the procedure in [7] and apply FedAVG on the online encoder and projector. We also adapt FedEMA [27], a notable federated SSL method originally developed to address the non-IID problem, as another method for performance comparison. FedEMA integrates BYOL as an SSL technique into its structure. Additional details regarding SSL baselines are presented in Appendix C. It is important to note that we avoid direct comparisons with FedDG [6], FedSR [9], FedADG [8], and FedDG-GA [24] since unlike our method, they employ the label information in their solutions. (2) Given the absence of prior research specifically addressing the problem of federated unsupervised domain generalization, we also compare FedGaLA to established centralized unsupervised domain generalization methods on the PACS and DomainNet datasets. We could not identify unsupervised domain generalization methods for Office-Home and TerraInc datasets. This comparison includes the following solutions: SimCLR [36], MoCo [50], BYOL [51], AdCo [52], and DARLING [45]. It is important to note that these methods are implemented in a non-federated environment and do not incorporate any data privacy constraints. All the results for these models are reported from [45].
@@
Implementation details. We use SimCLR as the SSL module in FedGaLA due to its performance on domain generalization problems as previously shown [45]. Following[45], ResNet-18 [53] is employed as the encoder network architecture for all experiments, which we train from scratch. We present details regarding data augmentations, projector architecture, and encoder hyperparameters in Appendix D. Following [47, 48], we first learn a representation by FedGaLA and the baseline models for 100 communication rounds with 7 local epochs. Next, we freeze the backbone model and train a liner classier for 100 epochs to perform prediction on the target domain. Appendix D presents all the hyperparameters used for FedGaLA. For FedEMA, we use the hyperparameters reported in [27]. All experiments were implemented using PyTorch and trained on 8 NVIDIA GeForce RTX 3090 GPUs.

### v2

@@ 3.3 FedGaLA
The analysis above establishes a link between gradient alignment and domain shift, forming the basis for our proposed FedGaLA framework. Figure 1 illustrates an overview of our framework for addressing the newly proposed problem setup (Definition 1). FedGaLA is remotely inspired by prior works that demonstrate improved generalization in centralized (non-federated) learning through gradient alignment (Mansilla et al. 2021; Shi et al. 2022; Rame, Dancette, and Cord 2022). However, our framework extends this notion by integrating gradient alignment into the field of unsupervised and federated learning, and assumes, unlike (Mansilla et al. 2021; Shi et al. 2022; Rame, Dancette, and Cord 2022), that the distribution of data from different clients is not known. Our core idea includes (i) enabling clients to learn domain-invariant representations at the client level through local gradient alignment, and (ii) adjusting the aggregation weights at the server level using global gradient alignment.
Algorithm 1 FedGaLA
1:Input: data $\mathcal{D}_{i}$, initialization $\Theta^{0}$
2:Output: $\Theta^{T}$
[…]
4:     Server:
6:           Initialize global update: $\textbf{g}^{(t+1)}\leftarrow\text{FedAVG}(\textbf{g}_{i}^{(t)})$
[…]
9:               Normalize weights: $w_{i}\leftarrow\frac{w_{i}}{\sum w_{i}}$
10:               Aggregate updates: $\hat{\textbf{g}}^{(t+1)}\leftarrow\sum_{i=1}^{K}w_{i}\textbf{g}_{i}^{(t)}$
11:     end for
13:           Communicate: $\Theta_{i}\leftarrow\Theta^{(t+1)}$
14:     Client:
[…]
20:                  Update weights:
21:                  $\theta_{i}^{(l,t)}\leftarrow\theta_{i}^{(l,t-1)}-\eta\cdot\nabla l_{i}^{(l,t)}(x_{i,j},\theta_{i}^{(l,t)})$
22:              end if
23:         end for
24:     end for
25:           Communicate: $S\leftarrow\Theta_{i}$
26:end for
At each communication round, clients are initialized with the global model. Subsequently, each client updates its parameters using SSL for $E$ epochs based on the local data through local gradient alignment and sends these updates back to the server. Finally, the server employs the global gradient alignment technique to perform aggregation. This procedure is repeated for $T$ communications rounds to determine the global model. The remaining part of this section provides details on local and global gradient alignment, and the complete framework is outlined in Algorithm 1.
@@
Local gradient alignment. Our method performs layer-wise local gradient alignment using a reference gradient. This reference is derived from the $l^{\text{th}}$ layer of the global model’s parameter updates between the current and the previous communication round. Suppose $\Theta=\{\theta^{(l)}\}_{l=1}^{L}$ indicates the parameters of the global model, where $\theta^{(l)}$ represents the parameters of the $l^{\text{th}}$ layer. The reference gradient for the $l^{th}$ layer is computed as $\hat{\textbf{g}}_{est}^{(l,t)}=\theta^{(l,t)}-\theta^{(l,t-1)}$, where $\theta^{(l,t)}$ and $\theta^{(l,t-1)}$ are the parameters of the $l^{\text{th}}$ layer of the global model at rounds $t$ and $t-1$, respectively. Then, $\hat{\textbf{g}}_{est}^{(l,t)}$ is locally used to determine whether the gradient of each layer obtained during training (e.g., via SGD) is aligned with the reference. The cosine similarity between the batch gradient and the reference for each layer $l$ at round $t$ is computed as $cos(\phi)^{(l,t)}=\frac{\langle\textbf{g}_{i,k}^{(l,t)},\textbf{g}_{\text{est}}^{(l,t)}\rangle}{\|\textbf{g}_{i,k}^{(l,t)}\|\cdot\|\textbf{g}_{\text{est}}^{(l,t)}\|}$, where $\textbf{g}_{i,k}^{(l,t)}$ is the gradient of $k^{\text{th}}$ batch of the $i^{\text{th}}$ client at layer $l$ and round $t$. Finally, batch gradients whose similarity with $\hat{\textbf{g}}_{est}^{(l,t)}$ are less than a user-defined threshold $\tau$ are discarded during the update process. This prevents clients from learning domain-specific features by disregarding local gradients that are not aligned with the global model. The rationale for discarding unaligned gradients instead of applying soft weighting is that when a gradient vector is unaligned with $\hat{\textbf{g}}_{est}$, scaling does not change this alignment, as the cosine of the angle between them is independent of scale. To establish a theoretical basis for the proposed local alignment, we introduce the following.
Proposition 1.
Given two sets of gradient vectors $\textbf{\emph{g}}_{i}$ and $\textbf{\emph{g}}_{j}$, by removing the $K^{th}$ vector in $\textbf{\emph{g}}_{j}$ where $cos(\textbf{\emph{g}}_{j,K},\textbf{\emph{g}}_{est})<0$, the covariance of two sets increases. For more details on this Proposition and its proof, see Appendix A.8.
Theorem 1 highlighted how an increase in domain shift between clients $C_{i}$ and $C_{j}$ correlates with a decrease in the covariance of their gradient vectors, $\textbf{g}_{i}$ and $\textbf{g}_{j}$. Therefore, gradient covariance can be used as an indicator of domain shift. Proposition 1 complements this by showing that by selectively removing gradients from $\textbf{g}_{j}$ that have a negative cosine similarity with an estimated target direction, $\textbf{g}_{est}$, we can effectively increase the covariance of the gradient sets, thus potentially counteracting the effects of domain shift.
@@
Global gradient alignment. To aggregate the local (client) models at the server, the locally measured gradients that closely match the average gradient across all clients are assigned greater weights. This soft weighting process operates as follows. Once the server receives the local models, it first obtains the local update $\hat{\textbf{g}}_{i}^{(t)}=\Theta_{i}^{(t)}-\Theta^{(t-1)}$. Then, the initial global update $\hat{\textbf{g}}^{(t+1)}$ is calculated by averaging all $\hat{\textbf{g}}_{i}^{(t)}$. Subsequently, the weight of each client is computed using $w_{i}=\frac{cos(\hat{\textbf{g}}_{i}^{(t)},\hat{\textbf{g}}^{(t+1)})+1}{2}$, where $w_{i}$ is the weight for the $i^{\text{th}}$ client and $\hat{\textbf{g}}_{i}^{(t)}$ is the gradient of the $i^{\text{th}}$ client at round $t$. This weight reflects the degree of alignment between each client’s update and the global model’s update direction. To ensure these weights are properly normalized using $w_{i}={w_{i}}/{\sum_{k=1}^{K}w_{k}}$. The normalization of weights allows for the proportional contribution of each client’s update.
@@
Finally, the normalized weights are used to perform aggregation. Each client’s model update is scaled by its respective weight, and these weighted updates are then aggregated to compute the weighted average update. The global model at round $t+1$ is updated based on $\hat{\textbf{g}}^{(t+1)}=\sum_{i=1}^{K}w_{i}\hat{\textbf{g}}_{i}^{(t)}$. This aggregation step is repeated three times, refining the weights with each iteration. This ensures the global model update is significantly influenced by clients whose updates align closely with the global learning objective. After completing the aggregation process, the global model is further updated based on $\Theta^{(t+1)}\leftarrow\Theta^{(t)}+\hat{\textbf{g}}^{(t+1)}$. The rationale for using different alignment strategies at the client and server levels stems from the inherent differences between local training and global aggregation. At the client level, we manage batch gradients, allowing us to specifically discard the unaligned ones without significant information loss. However, discarding gradients at the server-level corresponds to the deletion of entire clients, which can adversely affect the outcome.
@@ 4 Experiment setup
Datasets. To evaluate the effectiveness of our proposed method, we conduct experiments across four commonly used benchmarks for domain generalization. They include: PACS (Li et al. 2017), which consists of 9,991 images from four domains: ‘Photo’, ‘Art-painting’, ‘Cartoon’, and ‘Sketch’, across seven classes; Office-Home (Venkateswara et al. 2017), which consists of 15,588 images from four domains: ‘Art’, ‘Clipart’, ‘Product’, and ‘Real-world’, across 65 classes; TerraInc (Beery, Van Horn, and Perona 2018), which includes 24,788 images from four domains ‘Location 38’, ‘Location 43’, ‘Location 46’, and ‘Location 100’, across nine classes; DomainNet (Peng et al. 2019), which consists of 569,010 images in six domains: ‘Clipart’, ‘Infograph’, ‘Painting’, ‘Quickdraw’, ‘Real’, and ‘Sketch’, covering 345 classes. Following (Zhang et al. 2022), for the DomainNet dataset, we select the following classes: zigzag, tiger, tornado, flower, giraffe, toaster, hexagon, watermelon, grass, hamburger, blueberry, violin, fish, sun, broccoli, Eiffel tower, horse, train, bird, and bee (Zhang et al. 2022), resulting in a total of 38556 samples. In all experiments using this dataset, three domains are used for training (‘Painting’, ‘Real’, and ‘Sketch’) and the other three domains (‘Clipart’, ‘Infographics’, and ‘Quickdraw’) are used for testing.
Evaluation. We use the leave-one-domain-out setting used in prior works (Zhang et al. 2023a, 2022; Gulrajani and Lopez-Paz 2021). This involves selecting one domain as the target, training the model on the rest of the domains, and then testing the model’s performance on the selected target domain. Linear evaluation, a common feature evaluation approach, is utilized to evaluate the quality of learned representations (Feng, Xu, and Tao 2019; Zhang, Isola, and Efros 2017; Kolesnikov, Zhai, and Beyer 2019). For linear evaluation, following (van Berlo, Saeed, and Ozcelebi 2020; Zhuang, Wen, and Zhang 2022), we utilize 10% and 30% of the target data to train the linear classifier and evaluate the remaining 90% and 70% of the data, respectively.
Model Labeled PACS DomainNet
ratio P A C S Ave. C I Q Ave.
FedEMA 10% 50.0(0.7) 29.5(1.9) 42.4(2.3) 45.6(0.16) 41.9 38.6(1.1) 13.7(0.5) 45.5(1.6) 32.4
FedBYOL 52.1(1.1) 31.8(1.1) 45.4(2.2) 47.4(1.9) 44.2 38.1(0.5) 14.1(0.4) 53.6(2.9) 31.8
FedMoCo 58.5(2.2) 35.7(9.6) 37.7(12.9) 36.6(8.2) 42.1 30.5(0.6) 10.9(2.1) 46.4(0.8) 27.2
FedSimSiam 46.2(1.1) 28.6(1.0) 46.7(0.6) 37.6(1.3) 39.8 44.8(1.5) 12.2(0.3) 40.3(2.4) 36.9
FedSimCLR 64.2(1.2) 41.9(1.5) 58.4(1.3) 70.1(1.2) 58.6 45.2(0.4) 13.7(0.3) 59.7(0.7) 39.5
FedGaLA (ours) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0 47.6(0.9) 14.2(0.5) 61.4(0.3) 41.1
FedEMA 30% 53.5(0.3) 33.9(2.8) 48.3(2.4) 45.3(2.3) 45.2 43.5(0.8) 20.0(1.5) 49.0(2.8) 37.5
FedBYOL 55.1(1.2) 35.3(1.5) 48.3(1.5) 48.9(0.5) 46.9 44.6(1.2) 20.5(1.1) 50.4(0.9) 38.5
FedSimSiam 58.5(2.2) 35.7(9.6) 37.7(12.9) 36.6(8.2) 42.1 35.7(1.1) 14.9(3.9) 44.6(3.5) 31.8
FedMoCo 47.3(1.6) 30.2(3.4) 47.4(0.7) 34.4(1.7) 39.9 51.5(1.8) 18.2(0.1) 59.9(3.6) 43.2
FedSimCLR 69.8(1.1) 46.4(2.1) 63.9(1.6) 73.4(3.0) 63.3 51.7(1.0) 16.3(0.2) 66.5(0.9) 44.8
FedGaLA (ours) 71.1(2.0) 46.8(2.1) 65.7(1.6) 74.5(1.1) 64.6 52.4(0.7) 16.2(0.6) 68.8(0.6) 45.8
@@
Baselines. To evaluate our method, we take a two-pronged approach: (1) We adapt several popular SSL approaches to the federated domain generalization task, denoting them as FedSimCLR, FedMoCo, FedBYOL, and FedSimSiam. To this end, we employed SimCLR (Chen et al. 2020), MoCo (He et al. 2020), BYOL (Grill et al. 2020), and SimSiam (Chen and He 2021) in a federated setup. Each of these models consists of two encoders. BYOL and MoCo utilize exponential moving averages to update one of the encoders (target network) using the other encoder (online network). In contrast, SimCLR and SimSiam share weights between their two encoders. Additionally, SimCLR and MoCo, being contrastive-based SSL models, leverage negative samples in the learning process. Conversely, BYOL and SimSiam, which do not rely on negative samples, include a predictor on top of the online encoder to enhance learning from unlabeled samples. We then train each client locally using the respective SSL method. Next, we aggregate the trained encoders at the server using FedAVG (McMahan et al. 2017). For BYOL and SimSiam, we follow the procedure in (Zhuang et al. 2021) and apply FedAVG on the online encoder and projector. We also adapt FedEMA (Zhuang, Wen, and Zhang 2022), which is a commonly used method originally developed for federated unsupervised learning. FedEMA integrates BYOL as an SSL technique into its structure.
@@
It is important to note that we avoid direct comparisons with FedDG (Liu et al. 2021), FedSR (Nguyen, Torr, and Lim 2022), FedADG (Zhang et al. 2021b), FedIIR (Guo et al. 2023), and FedDG-GA (Zhang et al. 2023c) since unlike our method, they employ the label information in their solutions. (2) Given the absence of prior research specifically addressing the problem of federated unsupervised domain generalization, we also compare FedGaLA to established centralized unsupervised domain generalization methods on the PACS and DomainNet datasets. We could not identify unsupervised domain generalization methods for Office-Home and TerraInc datasets. This comparison includes the following solutions: SimCLR (Chen et al. 2020), MoCo (He et al. 2020), BYOL (Grill et al. 2020), AdCo (Hu et al. 2021), and DARLING (Zhang et al. 2022). It is important to note that these methods are implemented in a non-federated environment and do not incorporate any data privacy constraints. All the results for these models are reported from (Zhang et al. 2022).
@@
Implementation details. We use SimCLR as the SSL module in FedGaLA due to its performance on domain generalization problems as previously shown (Zhang et al. 2022). Following(Zhang et al. 2022), ResNet-18 (He et al. 2016) is employed as the encoder network architecture for all experiments, which we train from scratch. We present details regarding data augmentations, projector architecture, and encoder hyperparameters in Appendix B. Following (Feng, Xu, and Tao 2019; Zhang, Isola, and Efros 2017), we first learn a representation by FedGaLA and the baseline models for 100 communication rounds with 7 local epochs. Next, we freeze the backbone model and train a liner classier for 100 epochs to perform prediction on the target domain. Appendix B presents all the hyperparameters used for FedGaLA. For FedEMA, we use the hyperparameters reported in (Zhuang, Wen, and Zhang 2022). All experiments were implemented using PyTorch and trained on 8 NVIDIA GeForce RTX 3090 GPUs. For each experiment, we train the models three times with random initialization seeds and report the average.
Cross-silo vs. cross-device federated learning. Two frameworks can generally be considered in federated learning: cross-silo and cross-device. In cross-silo, the number of clients $K$ is small, and each client is capable of retaining its local model and state across training rounds. In contrast, cross-device federated learning works under the assumption of having millions of stateless clients, where only a limited number of clients participate in training during each communication round. However, due to limitations in experimental settings, the majority of studies typically conduct experiments involving at most hundreds of clients (Wang et al. 2020; Zhuang, Wen, and Zhang 2022). In this study, given the limited number of domains FedGaLA works within the cross-silo framework. Moreover, each client is equipped with the necessary resources to participate in the collaborative learning process. Additionally, we ensure that every client updates its local model with a sufficient number of epochs.
Figure 2: (a) Covariance under domain shift across various domains of the PACS dataset. (b) Average ratio of discarded local gradients over 100 communication rounds ($\tau=0$ and $E=1$). (c) Impact of various batch sizes. (d) Impact of $\tau$ on performance across different number of local epochs. Results in (b-d) are reported for the domain $P$ of the PACS dataset, with a 10% label ratio.
Model PACS DomainNet
P A C S Ave. C I Q Ave.
BYOL 27.0 25.9 21.0 19.7 23.4 14.6 8.7 5.9 9.7
MoCo 44.2 25.9 33.5 25.0 32.1 32.5 18.5 8.1 19.7
SimCLR 54.7 37.7 46.0 28.3 41.6 37.1 19.9 12.3 23.1
AdCo 46.5 30.2 31.5 22.9 32.8 32.3 17.9 11.6 20.6
DARLING 53.4 39.9 46.4 30.2 42.5 35.2 20.9 15.7 23.9
FedGaLA (ours) 64.7 44.2 60.5 70.5 60.0 47.6 14.2 61.4 41.1